OpenStudy (anonymous):

Find the roots of the polynomial equation. 2x3 + 2x2 – 19x + 20 = 0 I seriously don't understand this at all

OpenStudy (anonymous):

ok.. we start with the "rational root test" to find the first one

OpenStudy (anonymous):

p: factors of constant term 20 \(\pm1,\pm2,\pm4,\pm5,\pm10,\pm20\) q: factors of the coefficient of the highest power of "x" = 2 \(\pm1,\pm2\)

OpenStudy (anonymous):

then, by the rational-root theorem, if the given polynomial has a rational root, it must be one of the possible (p/q)\[ \pm{1\over1}=1,\pm{2\over1}=2,\pm{4\over1}=4,\ldots \]

OpenStudy (anonymous):

follow?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

evaluate

OpenStudy (anonymous):

ok, then by remainder theorm, we check each possibility one -by-one 1) x=1 \[2(1)^3+2(1)^2-19(1)+20\ne0\] so, x=1 is not a root 2) x=-1 \[2(-1)^3+2(-1)^2-19(-1)+20\ne0\] again, x=-1 is not a root

OpenStudy (anonymous):

3) x=2 \[2(2)^3+2(2)^2-19(2)+20\ne0\] not a root 4)x=-2 \[2(-2)^3+2(-2)^2-19(-2)+20\ne0\] not a root!!

OpenStudy (anonymous):

5) x=4 \[2(4)^3+2(4)^2-19(4)+20\ne0\] not a root 6)x=-4 \[2(-4)^3+2(-4)^2-19(-4)+20=\mathbf{0}\] Houston,m we have a winner

OpenStudy (anonymous):

we found one rational root. \(\Large{x = -4}\) for the next step, did you study the 'Descartes Rule of Signs'??

OpenStudy (anonymous):

don't know what that is

OpenStudy (anonymous):

not a problem.. but kapeesh so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good. Now, since the polynomial is of degree "3", there are '3' roots for it. and we found one for the other two, we have to take out what we found already.. so, \[(x+4)(\ldots)=2x^3+2x^2-19x+20\]

OpenStudy (anonymous):

the term in the paranthesis will be a quadratic (3-1=2) which we can get using polynomial division or sythethetic division

OpenStudy (anonymous):

you with me?

OpenStudy (anonymous):